Solve each system of equations using the substitution method.
4x - y = 32y = -2 x + 70

Answer by mananth(16946)

4x-y=32 .............1

y=70-2x
Plug the value of y in (1)
4x-1(70 -2x)=32
4x-70+2x=32
4x+2x=32 -70
6 x = 102
/ 6
x= 17
Plug the value of x in (1)
4 x + 1 y = 32
4 * 17 + 1 y = 32
68 + 1 y = 32
1 y = -36
/ 1
y= -36